# Unfitted finite element interpolated neural networks

### v1

@@ Abstract.
We present a novel approach that integrates unfitted finite element methods and neural networks to approximate partial differential equations on complex geometries. Easy-to-generate background meshes (e.g., a simple Cartesian mesh) that cut the domain boundary (i.e., they do not conform to it) are used to build suitable trial and test finite element spaces. The method seeks a neural network that, when interpolated onto the trial space, minimises a discrete dual norm (defined over the test space) of the weak residual functional associated to the equation. As with unfitted finite elements, essential boundary conditions are weakly imposed by Nitsche’s method. The method is robust to variations in Nitsche coefficient values, and to small cut cells. We experimentally demonstrate the method’s effectiveness in solving both forward and inverse problems across various 2D and 3D complex geometries, including those defined by implicit level-set functions and explicit stereolithography meshes. For forward problems with smooth analytical solutions, the trained neural networks achieve several orders of magnitude smaller $H^{1}$ errors compared to their interpolation counterparts. These interpolations also maintain expected $h$- and $p$-convergence rates. Moreover, using preconditioners (achieved by stabilisations) remarkably accelerates neural network training and further enhances robustness to the choice of Nitsche coefficient values. The experiments also show the method’s high accuracy and reliability in solving inverse problems, even with incomplete observations.
@@ 1. Introduction
The unfitted FEINN method transforms the (non)linear system of equations into a residual minimisation problem. The method is insensitive to small cut cell issues when residual vector-based loss functions (that is, based on algebraic norms applied to the residual vector) are employed. Even with loss functions based on the dual norm of the residual functional, only the linear test space requires stabilisation, which is a simpler task compared to stabilising the high-order trial space. A comprehensive set of experiments on 2D and 3D complex geometries demonstrates the effectiveness of unfitted FEINNs. We evaluate the method on linear and nonlinear forward PDEs, as well as inverse nonlinear problems, to confirm its reliability for solving PDEs with weak $H^{1}$ stability. The geometries considered in the experiments include implicitly defined ones via level-set functions, as well as those defined explicitly by stereolithography (STL) meshes, which are more representative of real-world applications [Badia2022b]. The trained NNs significantly outperform FE interpolations of the analytic solution in terms of $H^{1}$ error, particularly when the solution is smooth. Furthermore, when preconditioning is applied, we achieve further gains in accuracy, and the number of iterations required for convergence is substantially reduced.
@@ 2.6. Unfitted finite element interpolated neural networks
Let us now discuss the choice of the trial and test FE spaces. Since the FE problem (3) is reformulated as a minimisation problem (6), and the trial FE space is “hidden” in the PDE residual (4), we can choose the trial space to be $U_{h}^{\rm std}$, i.e., the standard unfitted FE space without any stabilisation. It is not necessary to apply stabilisation because the discrete minimisation problem is determined by the NN parameters, not the ill-posed DoFs in the FE space. In fact, all experiments in Sect. 4 utilise $U_{h}^{\rm std}$ as the trial space and yield accurate results. When NNs are interpolated onto $U_{h}^{\rm ag}$, they lose control over the cut cells during training because these DoFs are defined as linear combinations of the DoFs in the root cells. For the test FE space in the residual $\mathcal{R}_{h}$ (4), the choice depends on the specific loss function employed. Options include $V_{h}^{\rm std}$, with or without a ghost penalty, or $V_{h}^{\rm ag}$. Details on the loss functions and their corresponding test space selections are provided in Sect. 2.7. Moreover, as shown in [Badia2024, Badia2024adaptive], the linearisation of test spaces significantly accelerates convergence and improves the accuracy of NN solutions. In unfitted FEINNs, we also employ such Petrov-Galerkin FE formulation, adopting linearised $V_{h}^{\rm std}$ and $V_{h}^{\rm ag}$. These spaces are constructed by refining the background mesh discretising $\Omega^{\rm box}$ and building linear FE spaces on the active mesh resulting from intersecting the refined mesh and the domain boundary. Implementation details of these linearised unfitted test spaces are discussed in Sect. 3.1.
@@ 2.7. Loss functions
where $\left\lVert\cdot\right\rVert_{\chi}$ is an algebraic norm, e.g., $\ell^{1}$ or $\ell^{2}$ norm. Although the $\ell^{2}$ norm is more common in the literature [Berrone2022], the $\ell^{1}$ norm is also effective and can produce remarkable results, as revealed in [Badia2024, Badia2024adaptive]. We indeed use both norms in the experiments of this work. It is important to note that we always choose $V_{h}^{\rm std}$ as the test space when loss (7) is used. In this setup, neither the trial nor test spaces are stabilised, nor is there a need for ghost penalty terms to stabilise the discrete variational FE formulation. Nevertheless, as demonstrated in the moving domain tests in Sect. 4.1.2, the unfitted FEINN method remains robust to small cut cells. In this setting, the problem is stated as a residual minimisation, where several (local) minima do not disrupt the algorithm. This contrasts sharply with unfitted FEMs, where the solution requires multiplying the inverse of the system matrix and the right-hand side vector, facing ill-conditioning issues, i.e., a lack of uniqueness in the singular limit.
However, as mentioned in [Badia2024, Badia2024adaptive], the loss (7) becomes ill-posed as $h\rightarrow 0$; it amounts to computing the $L^{2}$-norm (up to a constant that depends on the quasi-uniformity of the mesh) of the residual functional, which is only in $H^{-1}(\Omega)$. Besides, using these algebraic norms on the residual vector, optimal convergence rates have not been proven [Berrone2022]. It has motivated the use of dual norms in [Badia2024], i.e.,
@@
In FEINNs, this dual norm can be computed by introducing a discrete Riesz projector $\mathcal{B}_{h}^{-1}:V_{h}^{\prime}\to V_{h}$ such that
@@
Taking $X=U$, the loss functions in (8) and (9) are equivalent. Other choices of $X$ are possible, e.g., $X=L^{2}(\Omega)$. To reduce computational costs, $\mathcal{B}_{h}$ can be substituted with any spectrally equivalent approximation; for example, geometric multigrid (GMG) preconditioners are explored in [Badia2024]. When utilising the preconditioned loss function (9), the test FE space can be either $V_{h}^{\rm std}$ or $V_{h}^{\rm ag}$. However, adding the ghost penalty (5) is essential to stabilise the variational problem (9) in $V_{h}^{\rm std}$, ensuring its invertibility in inexact arithmetic regardless of the small cuts. In contrast, discretising (9) in $V_{h}^{\rm ag}$ is inherently stable [Badia2018]. Indeed, in [Verdugo19], it was shown that standard multigrid solvers for body-fitted FEM reach expected mesh-independent convergence rates and linear complexity when applied to the Gram matrix in $V_{h}^{\rm ag}$. We compare the performance of these two test spaces in the experiments in Sect. 4.1.4.
The analysis of FEINNs for the loss function in (8) can be found in [Badia2024adaptive, Sect. 3] for strong imposition of boundary conditions and coercive systems. The extension to Nitsche’s method is straightforward, with minor modifications. These results can be used in the case of unfitted FEINNs under the assumption of a discrete inf-sup condition between the trial and test spaces (see [Badia2024adaptive, Sect. 2.2]) combined with the enhanced stability properties of CutFEM and AgFEM discretisation [Burman2010, Badia2018].
@@ 3.1. Unfitted linearised test spaces
We immediately observe that $V_{h}^{\rm std}$ has fewer DoFs than $U_{h}^{\rm std}$ (the same applies to $V_{h}^{\rm ag}$). A Petrov-Galerkin FE discretisation of the problem at hand results in a rectangular discrete system. While this mismatch is problematic for FEM, as it prevents the application of solvers for algebraic systems of equations, we can still apply the unfitted FEINN method because such method recasts the problem into a residual minimisation problem, which can be solved using a gradient-based optimiser. Although residual minimisation is also available in FEM (e.g., [Guermond2004]), the (body-fitted) FEINN method can produce more accurate NN solutions than FEM when the solution is smooth, as shown in [Badia2024, Badia2024adaptive, Badia2024compatible]. The computation of the discrete dual norm of the residual involves the inversion of a Galerkin discretisation operator built using the test space.
@@ 3.2. Unfitted finite element interpolated neural networks
We reuse the FEINN implementation as detailed in [Badia2024, Sect. 5] for the unpreconditioned and inverse problem cases, and in [Badia2024adaptive, Sect. 4.3] for the preconditioned case. For completeness, we highlight the main Julia packages used in the implementation of unfitted FEINNs. The FEM computations are managed by Gridap.jl [Badia2020, Verdugo2022], with unfitted extensions handled by GridapEmbedded.jl [GridapEmbeddedRepo] and STLCutters.jl [Martorell2024] packages. While GridapEmbedded.jl offers mesh intersection algorithms for implicit geometry representations, such as level set functions, STLCutters.jl handles explicit geometry representations using STL meshes. The neural network is implemented using Flux.jl [Innes2018]. We define custom reverse-mode rules in ChainRules.jl [ChainRules2024] for gradient computation.
@@ 4. Numerical experiments
We employ the same NN structure as in [Badia2024] for all the forward experiments, i.e., $L=5$ layers, $n=50$ neurons per hidden layer, and tanh activation. We use the BFGS optimiser from Optim.jl [Optimjl2018] to train NNs, following the settings in [Badia2024, Badia2024adaptive]. When the number of DoFs is relatively small (e.g., fewer than 1,000), NNs can easily overfit, so we reduce the number of training iterations (e.g., to a few hundreds). However, with sufficient number of DoFs, longer training does not negatively affect the training results, so we typically set a large number (e.g., 10,000) of iterations to ensure the NNs training process has converged. We also note that higher trial space orders require more iterations for the training to converge. Besides, we utilise the residual vector-based loss (7) as the default option for the experiments, as it has been proven to be effective in previous studies [Badia2024, Badia2024adaptive]. The benefits of using discrete dual norms (i.e., preconditioned losses) are highlighted in Sect. 4.1.4.
@@
where $u$ is the true solution, $\left\lVert\cdot\right\rVert_{L^{2}(\Omega)}=\sqrt{\int_{\Omega}|\cdot|^{2}}$, and $\left\lVert\cdot\right\rVert_{H^{1}(\Omega)}=\sqrt{\int_{\Omega}|\cdot|^{2}+|\boldsymbol{\nabla}(\cdot)|^{2}}$. After training, where the interpolated NN (i.e., $\pi_{h}(u_{\mathcal{N}})$) is used to compute the loss and gradients, we obtain a NN solution (i.e., $u_{\mathcal{N}}$), $u^{id}$ can be either $\pi_{h}(u_{\mathcal{N}})$ or $u_{\mathcal{N}}$. In the plots, these are labelled as “FEINN” and “NN”, respectively. For the $h$- and $p$-refinement error convergence tests, we repeat each experiment 10 times with different initial NN parameters to account for the sensitivity of achieved results with respect to NN parameter initialisation. Since the variations in the FEINN results are negligible, we only show the average errors among these 10 runs in the plots.
To establish a baseline to compare the accuracy of the identified solutions, we interpolate the true solution onto the trial FE space and compute the $L^{2}$ and $H^{1}$ errors. These interpolations, i.e., $\pi_{h}(u)$, are labelled as “FE interp” in the figures and their errors are referred to as $\left\lVert u-\pi_{h}(u)\right\rVert$ in the text. There are several reasons that prevent us from using FEM solutions as a baseline for comparison. Firstly, the FEM solutions are sensitive to the choice of Nitsche coefficient $\gamma$. Secondly, unfitted FEMs often require additional stabilisation techniques, such as ghost penalties in CutFEM, which introduce extra parameters that can affect solution accuracy. Lastly, as mentioned in Sect. 3.1, the Petrov-Galerkin FE formulation proposed in this work leads to a rectangular system that cannot be directly solved using standard sparse direct solvers. In contrast, computing the FE interpolation of the true solution is straightforward and unaffected by any hyperparameters.
@@
As shown in Fig. 2, we consider two 2D geometries: a disk and a flower. The disk’s level-set function is $\phi(x,y)=\sqrt{(x-x_{0})^{2}+(y-y_{0})^{2}}-R$, where $(x_{0},y_{0})$ is the centre and $R$ is the radius. The flower geometry, adapted from [Neiva2021], is described by $\phi(\rho,\theta)=\rho-0.12(\sin(5\theta)+3)$, where $\rho=\sqrt{(x-x_{0})^{2}+(y-y_{0})^{2}}$ and $\theta$ is the polar angle relative to the centre $(x_{0},y_{0})$. We use the flower shape in Sect. 4.1.5 and the disk shape in the remaining subsections. Unless otherwise specified, we employ the $\ell^{2}$ norm in the loss (7).

### v2

@@ Abstract.
We present a novel approach that integrates unfitted finite element methods and neural networks to approximate partial differential equations on complex geometries. Easy-to-generate background meshes (e.g., a simple Cartesian mesh) that cut the domain boundary (i.e., they do not conform to it) are used to build suitable trial and test finite element spaces. The method seeks a neural network that, when interpolated onto the trial space, minimises a discrete norm of the weak residual functional on the test space associated to the equation. As with unfitted finite elements, essential boundary conditions are weakly imposed by Nitsche’s method. The method is robust to variations in Nitsche coefficient values, and to small cut cells. We experimentally demonstrate the method’s effectiveness in solving both forward and inverse problems across various 2D and 3D complex geometries, including those defined by implicit level-set functions and explicit stereolithography meshes. For forward problems with smooth analytical solutions, the trained neural networks achieve several orders of magnitude smaller $H^{1}$ errors compared to their interpolation counterparts. These interpolations also maintain expected $h$- and $p$-convergence rates. Using the same amount of training points, the method is faster than standard PINNs (on both GPU and CPU architectures) while achieving similar or superior accuracy. Moreover, using a discrete dual norm of the residual (achieved by cut cell stabilisation) remarkably accelerates neural network training and further enhances robustness to the choice of Nitsche coefficient values. The experiments also show the method’s high accuracy and reliability in solving inverse problems, even with incomplete observations.
@@ 1. Introduction
The unfitted FEINN method transforms the (non)linear system of equations into a residual minimisation problem. The method is insensitive to small cut cell issues when residual vector-based loss functions (that is, based on algebraic norms applied to the residual vector) are employed. Even with loss functions based on the dual norm of the residual functional, only the computation of the discrete dual norm of the residual, i.e., an approximation of the $H^{1}$ inner product for the linearised FE space, requires stabilisation; this is a simpler task compared to stabilising the original unfitted problem with a high-order trial space and a linearised test space. A comprehensive set of experiments on 2D and 3D complex geometries demonstrates the effectiveness of unfitted FEINNs. We evaluate the method on linear and nonlinear forward PDEs, as well as inverse nonlinear problems, to confirm its reliability for solving PDEs with weak $H^{1}$ stability. The geometries considered in the experiments include implicitly defined ones via level-set functions, as well as those defined explicitly by stereolithography (STL) meshes, which are more representative of real-world applications [Badia2022b]. The method is compared against standard PINNs on both CPU and GPU architectures. Employing the same training points, unfitted FEINNs reach comparable accuracy using fewer training iterations and reduced computational time. The trained NNs significantly outperform FE interpolations of the analytic solution in terms of $H^{1}$ error, particularly when the solution is smooth. Furthermore, when a discrete dual norm of the residual is applied, we achieve further gains in accuracy, and the number of iterations required for convergence is substantially reduced.
@@ 2.6. Unfitted finite element interpolated neural networks
Let us now discuss the choice of the trial and test FE spaces. Since the FE problem (3) is reformulated as a minimisation problem (6), and the trial FE space is “hidden” in the PDE residual (4), we can choose the trial space to be $U_{h}^{\rm std}$, i.e., the standard unfitted FE space without any stabilisation. It is not necessary to apply stabilisation because the discrete minimisation problem is determined by the NN parameters, not the ill-posed DoFs in the FE space. In fact, all experiments in Sect. 4 utilise $U_{h}^{\rm std}$ as the trial space and yield accurate results. When NNs are interpolated onto $U_{h}^{\rm ag}$, they lose control over the cut cells during training because these DoFs are defined as linear combinations of the DoFs in the root cells. For the test FE space in the residual $\mathcal{R}_{h}$ (4), the choice depends on the specific loss function employed. Options include $V_{h}^{\rm std}$, with or without a ghost penalty, or $V_{h}^{\rm ag}$. Details on the loss functions and their corresponding test space selections are provided in Sect. 2.7. Moreover, as shown in [Badia2024, Badia2025adaptive], the linearisation of test spaces significantly accelerates convergence and improves the accuracy of NN solutions. In unfitted FEINNs, we also employ such Petrov-Galerkin FE formulation, adopting linearised $V_{h}^{\rm std}$ and $V_{h}^{\rm ag}$. These spaces are constructed by refining the background mesh discretising $\Omega^{\rm box}$ and building linear FE spaces on the active mesh resulting from intersecting the refined mesh and the domain boundary. Implementation details of these linearised unfitted test spaces are discussed in Sect. 3.1.
@@ 2.7. Loss functions
where $\left\lVert\cdot\right\rVert_{\chi}$ is an algebraic norm, e.g., $\ell^{1}$ or $\ell^{2}$ norm. When using a body-fitted mesh and the $\ell^{2}$ norm, the method becomes equivalent to the Nitsche variant of IVPINNs [Berrone2023]. However, in contrast to [Berrone2023], in this work we focus on unfitted meshes to handle complex geometries. When the loss (7) is employed for unfitted FEINNs, the method is referred to as $\ell^{1}$- or $\ell^{2}$-FEINNs, depending on the algebraic norm chosen. Although the $\ell^{2}$ norm is more common in the literature [Berrone2022], the $\ell^{1}$ norm is also effective and can produce remarkable results, as revealed in [Badia2024, Badia2025adaptive]. We indeed use both norms in the experiments of this work. It is important to note that we always choose $V_{h}^{\rm std}$ as the test space when loss (7) is used. In this setup, neither the trial nor test spaces are stabilised, nor is there a need for ghost penalty terms to stabilise the discrete variational FE formulation. Nevertheless, as demonstrated in the moving domain tests in Sect. 4.1.3, the unfitted FEINN method remains robust to small cut cells. In this setting, the problem is stated as a residual minimisation, where several (local) minima do not disrupt the algorithm. This contrasts sharply with unfitted FEM, where the solution requires multiplying the inverse of the system matrix and the right-hand side vector, facing ill-conditioning issues, i.e., a lack of uniqueness in the singular limit.
However, as mentioned in [Badia2024, Badia2025adaptive], the loss (7) becomes ill-posed as $h\rightarrow 0$; it amounts to computing the $L^{2}$-norm (up to a constant that depends on the quasi-uniformity of the mesh) of the residual functional, which is only in $H^{-1}(\Omega)$. Besides, using these algebraic norms on the residual vector, optimal convergence rates have not been proven [Berrone2022]. It has motivated the use of dual norms in [Badia2024], i.e.,
@@
In FEINNs, this discrete dual norm can be computed by introducing a discrete Riesz projector $\mathcal{B}_{h}^{-1}:V_{h}^{\prime}\to V_{h}$ such that
@@
When $X=U$, the loss functions in (8) and (9) are equivalent, and the method is denoted as $H^{-1}$-FEINNs. Other choices of $X$ are possible, e.g., $X=L^{2}(\Omega)$. To reduce computational costs, $\mathcal{B}_{h}$ can be substituted with any spectrally equivalent approximation; for example, geometric multigrid (GMG) preconditioners are explored in [Badia2024].
When utilising the loss function with the discrete dual norm of the residual (9), the test FE space can be either $V_{h}^{\rm std}$ or $V_{h}^{\rm ag}$. However, adding the ghost penalty (5) is essential to stabilise the variational problem (9) in $V_{h}^{\rm std}$, ensuring its invertibility in inexact arithmetic regardless of the small cuts. In contrast, discretising (9) in $V_{h}^{\rm ag}$ is inherently stable [Badia2018]. Indeed, in [Verdugo19], it was shown that standard multigrid solvers for body-fitted FEM reach expected mesh-independent convergence rates and linear complexity when applied to the Gram matrix in $V_{h}^{\rm ag}$. We compare the performance of these two test spaces in the experiments in Sect. 4.1.5.
The analysis of FEINNs for the loss function in (8) can be found in [Badia2025adaptive, Sect. 3] for strong imposition of boundary conditions and coercive systems. The extension to Nitsche’s method is straightforward, with minor modifications. These results can be used in the case of unfitted FEINNs under the assumption of a discrete inf-sup condition between the trial and test spaces (see [Badia2025adaptive, Sect. 2.2]) combined with the enhanced stability properties of CutFEM and AgFEM discretisation [Burman2010, Badia2018].
@@ 3.1. Unfitted linearised test spaces
We immediately observe that $V_{h}^{\rm std}$ has fewer DoFs than $U_{h}^{\rm std}$ (the same applies to $V_{h}^{\rm ag}$). A Petrov-Galerkin FE discretisation of the problem at hand results in a rectangular discrete system. While this mismatch is problematic for FEM, as it prevents the application of solvers for algebraic systems of equations, we can still apply the unfitted FEINN method because such method recasts the problem into a residual minimisation problem, which can be solved using a gradient-based optimiser. Although residual minimisation is also available in FEM (e.g., [Guermond2004]), the (body-fitted) FEINN method can produce more accurate NN solutions than FEM when the solution is smooth, as shown in [Badia2024, Badia2025adaptive, Badia2024compatible]. The computation of the discrete dual norm of the residual involves the inversion of a Galerkin discretisation operator built using the test space.
@@ 3.2. Unfitted finite element interpolated neural networks
We reuse the FEINN implementation as detailed in [Badia2024, Sect. 5] for the Euclidean norm of the residual as loss function and inverse problem cases, and in [Badia2025adaptive, Sect. 4.3] when using a discrete dual norm of the residual as loss function. For completeness, we highlight the main Julia packages used in the implementation of unfitted FEINNs. The FEM computations are managed by Gridap.jl [Badia2020, Verdugo2022], with unfitted extensions handled by GridapEmbedded.jl [GridapEmbeddedRepo] and STLCutters.jl [Martorell2024] packages. While GridapEmbedded.jl offers mesh intersection algorithms for implicit geometry representations, such as level set functions, STLCutters.jl handles explicit geometry representations using STL meshes. The neural network is implemented using Flux.jl [Innes2018]. We define custom reverse-mode rules in ChainRules.jl [ChainRules2024] for gradient computation.
@@ 4. Numerical experiments
We employ the same NN structure as in [Badia2024] for all the forward experiments, i.e., $L=5$ layers, $n=50$ neurons per hidden layer, and tanh activation. Unless otherwise specified, we use the BFGS optimiser from Optim.jl [Optimjl2018] to train NNs, following the settings in [Badia2024, Badia2025adaptive]. When the number of DoFs is relatively small (e.g., fewer than 1,000), NNs can easily overfit, so we reduce the number of training iterations (e.g., to a few hundreds). However, with sufficient number of DoFs, longer training does not negatively affect the training results, so we typically set a large number (e.g., 10,000) of iterations to ensure the NNs training process has converged. We also note that higher trial space orders require more iterations for the training to converge. Besides, by default, we use $\ell^{2}$-FEINNs (i.e., the Euclidean norm of the residual in loss function (7)), which has been proven to be effective for smooth solutions [Badia2024, Badia2025adaptive]. For rough solutions with localised high gradients, we use $H^{-1}$-FEINNs, i.e., loss (9) with $X=H^{1}$, which notably improves robustness as shown in Sect. 4.1.5.
@@
where $u$ is the true solution, $\left\lVert\cdot\right\rVert_{L^{2}(\Omega)}=\sqrt{\int_{\Omega}|\cdot|^{2}}$, and $\left\lVert\cdot\right\rVert_{H^{1}(\Omega)}=\sqrt{\int_{\Omega}|\cdot|^{2}+|\boldsymbol{\nabla}(\cdot)|^{2}}$. After training, where the interpolated NN (i.e., $\pi_{h}(u_{\mathcal{N}})$) is used to compute the loss and gradients, we obtain a NN solution (i.e., $u_{\mathcal{N}}$), $u^{id}$ can be either $\pi_{h}(u_{\mathcal{N}})$ or $u_{\mathcal{N}}$. For the $h$- and $p$-refinement error convergence tests, we repeat each experiment 10 times with different initial NN parameters to account for the sensitivity of achieved results with respect to NN initialisation. Since the variations in the $\pi_{h}(u_{\mathcal{N}})$ solutions are negligible, we only show the average errors among these 10 runs in the plots.
To establish a baseline to compare the accuracy of the identified solutions, we interpolate the true solution onto the trial FE space and compute the $L^{2}$ and $H^{1}$ errors. These interpolations are denoted as $\pi_{h}(u)$ in the figures and their errors are referred to as $\left\lVert u-\pi_{h}(u)\right\rVert$ in the text. There are several reasons that prevent us from using FEM solutions as a baseline for comparison. Firstly, the FEM solutions are sensitive to the choice of Nitsche coefficient $\gamma$. Secondly, unfitted FEMs often require additional stabilisation techniques, such as ghost penalties in CutFEM, which introduce extra parameters that can affect solution accuracy. Lastly, as mentioned in Sect. 3.1, the Petrov-Galerkin FE formulation proposed in this work leads to a rectangular system that cannot be directly solved using standard sparse direct solvers. In contrast, computing the FE interpolation of the true solution is straightforward and unaffected by these hyperparameters.
@@
As shown in Fig. 2, we consider two 2D geometries: a disk and a flower. The disk’s level-set function is $\phi(x,y)=\sqrt{(x-x_{0})^{2}+(y-y_{0})^{2}}-R$, where $(x_{0},y_{0})$ is the centre and $R$ is the radius. The flower geometry, adapted from [Neiva2021], is described by $\phi(\rho,\theta)=\rho-0.12(\sin(5\theta)+3)$, where $\rho=\sqrt{(x-x_{0})^{2}+(y-y_{0})^{2}}$ and $\theta$ is the polar angle relative to the centre $(x_{0},y_{0})$. We use the flower shape in Sect. 4.1.6 and the disk shape in the remaining subsections.